# Finding Stable Price Zones in European Electricity Markets: Aiming to Square the Circle?

### v2

@@ 5.1 Analysis of the price zone configurations proposed by ACER
The price zone configurations proposed by ACER are not stable over time. Given the nodal prices corresponding to 1989, 1995 and 2009, respectively, K-Means and Spectral Clustering compute price zones that are considerably different from the proposed configurations.
We recall from Section 3.2 that the Rand index measures the similarity between two clusterings by considering all pairs of nodes and counting the pairs that are assigned to the same or to different clusters in both clusterings. As Table 1 indicates, Rand indices vary considerably when different years are analyzed. For instance, for DE4 (spectral) and the clustering computed with Spectral Clustering with price and location features we get $RI=0.56$ for 1989, $RI=0.46$ for 1995 and $RI=0.56$ for 2009. Thus, our analysis suggests that the configurations proposed by ACER are not stable over time, regardless of the number of price zones. An explanation is that nodal prices vary both temporally and spatially. As a result, the spatial distribution of the optimal clusters (i.e., obtained based on the prices corresponding to a specific time frame) is different each year. The plots included in Appendix D show that the clusterings indeed differ across time.
5.2 Analysis of the optimal price zone configurations for 1989, 1995 and 2009
In what follows, we analyze alternative price zones computed with K-Means and Spectral Clustering for the years 1989, 1995 and 2009. Previously we showed that the configurations proposed by ACER are not stable over these years. Now we are interested to see whether other price zone configurations exist that are more stable with regards to the criteria defined in Section 3.2. We report the results of K-Means and Spectral Clustering as described in Section 3.1. We evaluate the clusterings computed for each year using (1) time series of nodal prices as features, and (2) time series of nodal prices and the location coordinates of nodes.
@@ 5.2 Analysis of the optimal price zone configurations for 1989, 1995 and 2009
The clusterings computed with K-Means for 1989, 1995 and 2009 based only on prices exhibit lower average price standard deviations compared to the configurations proposed by ACER.
@@
We observe that the average price standard deviations $\sigma$ differ depending on the clustering algorithm. For instance, for $k=3$, K-Means (Spectral Clustering) leads to $\sigma=12.11(14.81)$ for 1989, $\sigma=12.47(15.56)$ for 1995 and $\sigma=10.06(11.11)$ for 2009. However, regardless of the year and the number of clusters, the clusterings we compute with K-Means based on nodal prices lead to average price standard deviations that are generally lower compared to the configurations proposed by ACER. Recall that the average price standard deviations for the configurations proposed by ACER are up to 13.98. Nevertheless, the clusters we compute with K-Means and Spectral Clustering are not necessarily balanced, that is, the clusters may not contain a similar number of nodes. This fact can be observed in the plots provided in Appendix D. Given this observation and the fact that the configurations proposed by ACER appear to be balanced (see Appendix A), it explains that the price standard deviations inside the clusters we compute with K-Means and Spectral Clustering are lower.
@@
The clusterings computed with K-Means and Spectral Clustering for 1989, 1995 and 2009 are characterized by low inter-cluster dissimilarity and are thus not well-defined. Regardless of algorithm or clustering features, the yearly differences between the average cluster prices are less than 6 EUR/MWh and 3.45 EUR/MWh on average.
@@
Given these observations, we conclude that the price zones computed for 1989, 1995 and 2009 with K-Means and Spectral Clustering generally exhibit larger intra-cluster similarity and similar inter-cluster dissimilarity compared to the configurations proposed by ACER. Large intra-cluster similarity indicates strong cohesion inside clusters and is a consequence of unbalanced clusters relative to the ACER configurations.
@@
The clusterings computed with K-Means and Spectral Clustering are not temporally stable. A signficant number of nodes would be assigned to different clusters in different years.
@@
Last but not least, we analyze how spatially coherent the clusterings computed with K-Means and Spectral Clustering for 1989, 1995 and 2009 are. Figure 2 illustrates the Global Moran’s I values of the clusterings computed using (1) nodal weekly prices and (2) nodal weekly prices and the location coordinates of nodes as features.
@@
Analyzing the clusters on the map of Germany, we notice that isolated points assigned to a specific cluster are sometimes located in the geographic area predominantly occupied by points belonging to a different cluster. This behavior can also be observed when latitude and longitude are included as features and is independent of the number of clusters or clustering algorithm. Therefore, a post-processing step is required to obtain spatially coherent price zones. This is because prices alone do not define geographically well-separated or spatially coherent clusters. This phenomenon can be observed in Figures 12 and 11 from Appendix D.
@@
6 Conclusions and Policy Implications
@@
In this paper, we analyzed the alternative price zone configurations proposed by ACER for Germany, namely DE2 (k-means), DE2 (spectral), DE3 (spectral) and DE4 (spectral). Similar to ACER’s approach (the algorithms used are publicly known, but the implementation details are not), we computed alternative price zone configurations by solving a clustering problem using K-Means and Spectral Clustering. Clusterings that we obtained with K-Means using only latitude and longitude as features are similar to the configurations proposed by ACER. Additionally, we computed alternative configurations by taking nodal prices or nodal prices and geographic coordinates into account and compared them with the configurations proposed by ACER. The resulting price zones were different from the proposed ACER configurations. Moreover, our results show that clusters with “minimal price dispersion within each bidding zone” (ACER, 2022a ) were not geographically coherent on the map, and they do not have similar numbers of nodes. Note that ACER might have had access to a more complete dataset covering all substations and also the clustering algorithms might differ in details.
@@
We evaluated how different configurations proposed by ACER reduce the average price standard deviations within clusters, and found that the configurations do not reduce the price standard deviations within zones much. Moreover, the decrease in average price standard deviation is not monotonic with respect to the number of clusters. Configurations with 3 or 4 price zones do not reduce the average price standard deviation as compared to configurations with 1 or 2 price zones. We also observed that the average prices of all proposed configurations do not differ considerably across clusters. For all three years we considered, the annual differences between the average cluster prices are less than 6 EUR/MWh.
Independent of how the clusterings were determined and whether we look at those determined by ACER or our own, our main finding is that none of them is stable across time. What appears to be a good cluster in one time period might turn out suboptimal in another time period.
Note that our analysis is based on data from the Bidding Zone review with a target year 2025. With the ambitious goals for renewable energy in Germany until 2030 and beyond, and the increasing electrification of the traffic and heat sectors, supply and demand will change substantially and the analysis might soon be outdated. Splitting price zones incurs high transition costs and there is a danger that new reconfigurations of the bidding zones might be in order in a few years again.
@@ B.1 Cluster price standard deviations
The following plots illustrate, for every configuration, the cluster price standard deviation corresponding to each week in 1989, 1995, and 2009. A point represents the price standard deviation of a specific cluster in a specific week. Depending on the configuration (i.e., number of clusters), multiple points are shown for the same week. Figure 5 highlights the weekly price standard deviation of a single price zone for each year we consider.
Figure 4: Single price zone (cluster)
@@ Appendix E Single-day clustering analysis
The time granularity of the prices included in the clustering computation influences the stability of the resulting clusterings. In what follows, we analyze the stability of the clusterings identified given the prices corresponding to a single day. We select the day of January 22, 1989, and compute clusterings using K-Means and Spectral Clustering. As clustering feature(s), we consider (1) the time series of prices of a node (12 hours, i.e., 12 features) and (2) the time series of prices of a node and its geographic coordinates.

### v4

@@ 5.1 Analysis of the price zone configurations proposed by ACER
The price zone configurations proposed by ACER are not stable over time. Given the nodal prices corresponding to the scenario years 1989, 1995 and 2009, respectively, K-Means and Spectral Clustering compute price zones that are considerably different from the proposed configurations.
We recall from Section 3.2 that the Rand index measures the similarity between two clusterings by considering all pairs of nodes and counting the pairs that are assigned to the same or different clusters in both clusterings. As Table 1 indicates, Rand indices vary considerably when different years are analyzed. For instance, for DE4 (spectral) and the clustering computed with Spectral Clustering with price and location features we get $RI=0.56$ for 1989 yr scenario, $RI=0.46$ for 1995 yr scenario and $RI=0.56$ for 2009 yr scenario. Thus, our analysis suggests that the configurations proposed by ACER are not stable over time, regardless of the number of price zones (see also Appendix D).
There are two main reasons for temporally unstable configurations. The first reason is that nodal prices vary both temporally and spatially due to various factors, such as supply and demand fluctuations, fuel prices and weather conditions. The second reason is related to the non-robustness of the clustering algorithms that we consider. K-Means is sensitive to the initial placement of centroids and outliers, which can significantly affect the resulting clusters. Spectral Clustering also exhibits sensitivity to the choice of parameters such as the implemented similarity metric.
5.2 Analysis of the optimal price zone configurations for the scenario years 1989, 1995 and 2009
In what follows, we analyze alternative price zones computed with K-Means and Spectral Clustering for the scenario years 1989, 1995 and 2009. Previously we showed that the configurations proposed by ACER are not stable over these years. Now we are interested to see whether other price zone configurations exist that are more stable with regards to the criteria defined in Section 3.2. We report the results of K-Means and Spectral Clustering as described in Section 3.1. We evaluate the clusterings computed for each year using (1) time series of nodal prices as features, and (2) time series of nodal prices and the location coordinates of nodes.
@@ 5.2 Analysis of the optimal price zone configurations for the scenario years 1989, 1995 and 2009
The clusterings computed with K-Means for the scenario years 1989, 1995 and 2009 based only on prices exhibit lower average price standard deviations compared to the configurations proposed by ACER.
@@
We observe that the average price standard deviations $\sigma$ differ depending on the clustering algorithm. For instance, for $k=3$, K-Means (Spectral Clustering) leads to $\sigma=12.11(14.81)$ for 1989 yr scenario, $\sigma=12.47(15.56)$ for 1995 yr scenario and $\sigma=10.06(11.11)$ for 2009 yr scenario. However, regardless of the year and the number of clusters, the clusterings we compute with K-Means based on nodal prices lead to average price standard deviations that are generally lower compared to the configurations proposed by ACER. Recall that the average price standard deviations for the configurations proposed by ACER are up to 13.98. Nevertheless, the clusters we compute with K-Means and Spectral Clustering are not necessarily balanced, that is, the clusters may not contain a similar number of nodes. This fact can be observed in the plots provided in Appendix D. Given this observation and the fact that the configurations proposed by ACER appear to be balanced (see Appendix A), it explains that the price standard deviations inside the clusters we compute with K-Means and Spectral Clustering are lower.
@@
The clusterings computed with K-Means and Spectral Clustering for the scenario years 1989, 1995 and 2009 are characterized by low inter-cluster dissimilarity and are thus not well-defined. Regardless of algorithm or clustering features, the yearly differences between the average cluster prices are less than 6 EUR/MWh and 3.45 EUR/MWh on average.
@@
Given these observations, we conclude that the price zones computed for the scenario years 1989, 1995 and 2009 with K-Means and Spectral Clustering generally exhibit larger intra-cluster similarity and similar inter-cluster dissimilarity compared to the configurations proposed by ACER. Large intra-cluster similarity indicates strong cohesion inside clusters and is a consequence of unbalanced clusters relative to the ACER configurations.
@@
The clusterings computed with K-Means and Spectral Clustering are not temporally stable. A significant number of nodes would be assigned to different clusters in different years.
@@
Last but not least, we analyze how spatially coherent the clusterings computed with K-Means and Spectral Clustering for the scenario years 1989, 1995 and 2009 are. Figure 4 illustrates the Global Moran’s I values of the clusterings computed using (1) nodal weekly prices and (2) nodal weekly prices and the location coordinates of nodes as features.
@@
Analyzing the clusters on the map of Germany, we notice that isolated points assigned to a specific cluster are sometimes located in the geographic area predominantly occupied by points belonging to a different cluster. This behavior can also be observed when latitude and longitude are included as features and is independent of the number of clusters or clustering algorithm. Therefore, a post-processing step is required to obtain spatially coherent price zones. This is because prices alone do not define geographically well-separated or spatially coherent clusters. This phenomenon can be observed in Figures 15 and 14 from Appendix D.
@@
5.3 Effects of unstable zonal configurations
Our experiments show that the alternative bidding zone configurations for Germany exhibit low price similarity within zones and high price similarity between zones, and are not stable over time. Unstable configurations with respect to these criteria have negative implications related to equity and market efficiency.
The density plots from Figure 5 illustrate the distribution of average nodal prices in the two zones of the DE2 (k-means) configuration, with the average zonal prices highlighted for comparison. The average nodal prices were computed considering all prices from the LMP Study (see Section 4). We observe that the distribution corresponding to zone 1 has long tails, which indicates large variations in nodal prices within zones. These variations lead to regional inequities and distort the zonal price signals, meaning the prices do not accurately reflect the supply and demand conditions in specific subregions. This results in poor local incentives to adapt to energy prices, poor investment incentives, and an inefficient implementation of scarcity pricing (Papavasiliou, 2020). Furthermore, the distribution for zone 2 reveals that over 50% of the nodes in this zone have average prices below the average zonal price. If the uniform price for this zone was set at the average zonal price, it would create equity issues among the market agents located at these nodes.
Figure 5: Density plots showing the distribution of average nodal prices (EUR/MWh) in the two zones of the DE2 (k-means) configuration, with the average zonal prices highlighted for comparison
Our analysis suggests that nodes with similar prices are assigned to different zones under the alternative configurations being discussed (see Appendix B). This discrepancy can create economic inefficiencies and fairness issues if consumers and producers in these zones experience different electricity prices (Morales & Hanly, 2018).
Morevoer, temporally unstable zonal configurations (Table 1) can lead to long-term market inefficiencies by distorting price signals, reducing market liquidity, and creating investment uncertainty (Gugler et al., 2020). Inconsistent price patterns over time can diminish market liquidity by introducing uncertainty and disrupting trading activities. Moreover, this uncertainty about future market conditions can discourage investment in new generation capacity and infrastructure upgrades.
6 Conclusions
@@
In this paper, we analyzed the alternative price zone configurations proposed by ACER for Germany, namely DE2 (k-means), DE2 (spectral), DE3 (spectral) and DE4 (spectral). Similar to ACER’s approach (the algorithms used are publicly known, but the implementation details are not), we computed alternative price zone configurations by solving a clustering problem using K-Means and Spectral Clustering. Clusterings that we obtained with K-Means using only latitude and longitude as features are similar to the configurations proposed by ACER. Additionally, we computed alternative configurations by taking nodal prices or nodal prices and geographic coordinates into account and compared them with the configurations proposed by ACER. The resulting price zones were different from the proposed ACER configurations. Moreover, our results show that clusters with “minimal price dispersion within each bidding zone” (ACER, 2022a ) were not geographically coherent on the map, and they do not have similar numbers of nodes. Note that ACER might have had access to a more complete dataset such that a one-to-one comparison to their results is not possible. We can only leverage the data that was made available publicly.
We evaluated the stability of the alternative configurations considering three criteria, namely intra- and inter-cluster similarity, temporal stability and spatial coherence. Our findings have important implications concerning the ongoing discussion of whether a split of the German bidding zone should be performed.
@@
Firstly, splitting the German bidding zone into 2, 3 or 4 zones does not reduce the price dispersion within zones by more than 3-9%. Moreover, the decrease in average price standard deviation is not monotonic with respect to the number of zones. Configurations with 3 or 4 zones do not necessarily reduce the average price standard deviation compared to configurations with fewer zones. Furthermore, for all proposed configurations, the average prices across zones do not differ considerably. The annual differences between the average cluster prices are less than 6 EUR/MWh for all three years we analyzed. As discussed in (ACER, 2020), two main goals of a split are to reduce price dispersion within individual bidding zones and to reflect long-term differences in average prices across the resulting zones. Our analysis shows that the alternative configurations proposed by ACER can be challenged based on these objectives.
Secondly, independent of how the alternative price zone configurations were determined and whether we look at those suggested by ACER or our own, none of them is temporally stable with respect to prices. Recall that prices were the only clustering feature in ACER’s methodology for obtaining alternative configurations (ACER, 2022b ) and “can be used as proxies for economic efficiency(ACER, 2022a ). Our study shows that a configuration that is optimal in one time period might turn out to be suboptimal in another. There can be many arguments for or against a bidding zone split. However, when a decision should be made regarding the configuration of Germany into bidding zones, the data revealed and the method defined during the BZR do not allow for a clear recommendation.
Data Availability
The data used for this publication are available on the Bidding Zone Review website (ENTSO-E, 2023).
@@ B.1 Cluster price standard deviations
The following plots illustrate, for every configuration, the cluster price standard deviation corresponding to each week in the scenario years 1989, 1995, and 2009. A point represents the price standard deviation of a specific cluster in a specific week. Depending on the configuration (i.e., number of clusters), multiple points are shown for the same week. Figure 8 highlights the weekly price standard deviation of a single price zone for each year we consider.
@@ Appendix E Single-day clustering analysis
The time granularity of the prices included in the clustering computation influences the stability of the resulting clusterings. In what follows, we analyze the stability of the clusterings identified given the prices corresponding to a single day. We select the scenario day of January 22, 1989, and compute clusterings using K-Means and Spectral Clustering. As clustering feature(s), we consider (1) the time series of prices of a node (12 hours, i.e., 12 features) and (2) the time series of prices of a node and its geographic coordinates.